ABC Company has set the following standards in order to produce one unit of its single product: standard quantity standard price direct materials 2.7 yards ?? per yard direct labor 2 hours $14 per hour variable overhead 2 hours ?? per hour During August, ABC Company spent $133,770 to purchase direct materials and had direct labor totaling $189,440. During August, ABC Company used 16,000 yards of direct materials in the production of units. ABC had 2,500 yards of materials in its August 1 direct material inventory and had 1,200 yards yards remaining in its direct materials inventory at August 31. ABC Company reported the following variances for August: Direct material price variance .............. $12,495 unfavorable Direct labor rate variance .................. $24,240 unfavorable Total direct labor variance ................. $6,560 favorable Variable overhead spending variance ......... $1,900 unfavorable Variable overhead efficiency variance ....... $20,900 favorable Calculate ABC's direct material quantity variance for August. If the variance is favorable, place a minus sign in front of your answer (i.e., -5000). If the variance is unfavorable, simply enter your answer as a number (i.e., 5000).
Answer #1

Total direct labor variance = $6,560 Favorable

Actual direct labor cost = $189,440

Standard direct labor cost = Actual direct labor cost +Favorable direct labor variance

= $189,440 + $6,560 = $196,000

Standard labor cost per unit = 2*$14 = $28 per unit

Nos of units produced by ABC company in august = Standard direct labor cost / Standard labor cost per unit = $196,000 / $28 = 7000 units

Material quantity purchased = 16000+1200-2500 = 14700 yards

Materials price variance = (Standard price - actual price) x Material purchased

- 12495 = (Standard Price per yard - 133770/14700)*14700

Standard Price per yard = $8.25 per unit

Materials quantity variance = (Standard quantity - actual quantity used) x standard price
= (7000*2.7 - 16000)*8.25

= $23,925 Favorable